### v2

# PASS: Probabilistic Agentic Supernet Sampling for Interpretable and Adaptive Chest X-Ray Reasoning

Abstract
Existing tool-augmented agentic systems are limited in the real world by (i) black-box reasoning steps that undermine trust of decision-making and pose safety risks, (ii) poor multimodal integration, which is inherently critical for healthcare tasks, and (iii) rigid and computationally inefficient agentic pipelines. We introduce PASS (Probabilistic Agentic Supernet Sampling), the first multimodal framework to address these challenges in the context of Chest X-Ray (CXR) reasoning. PASS adaptively samples agentic workflows over a multi-tool graph, yielding decision paths annotated with interpretable probabilities. Given the complex CXR reasoning task with multimodal medical data, PASS leverages its learned task-conditioned distribution over the agentic supernet. Thus, it adaptively selects the most suitable tool at each supernet layer, offering probability-annotated trajectories for post-hoc audits and directly enhancing medical AI safety. PASS also continuously compresses salient findings into an evolving personalized memory, while dynamically deciding whether to deepen its reasoning path or invoke an early exit for efficiency. To optimize a Pareto frontier balancing performance and cost, we design a novel three-stage training procedure, including expert knowledge warm-up, contrastive path-ranking, and cost-aware reinforcement learning. To facilitate rigorous evaluation, we introduce CAB-E, a comprehensive benchmark for multi-step, safety-critical, free-form CXR reasoning. Experiments across various benchmarks validate that PASS significantly outperforms strong baselines in multiple metrics (e.g., accuracy, AUC, LLM-J.) while balancing computational costs, pushing a new paradigm shift towards interpretable, adaptive, and multimodal medical agentic systems.
1 Introduction
Figure 1: An overview of PASS. Given a multimodal complex reasoning task (CXR image, textual comprehensive query, multimodal personalized context), our probabilistic controller learns a continuous task-conditioned distribution over the agentic supernet (i.e. a directed acyclic graph of medical agent containers). At each step, it samples an action, yielding a workflow annotated with interpretable probabilities for post-audits and directly enhances clinical AI safety . Tool outputs, which can be both text and images, are summarized and fed into an evolving personalized memory and shared in-context to inform subsequent steps. The controller is trained via a principled three-stage strategy (expert knowledge warm-up, contrastive path ranking, cost-aware reinforcement learning) to optimize the accuracy-cost trade-off. Eventually, PASS is enabled to answer multimodal medical questions in free-form text via an interpretable, adaptive, and efficient agentic reasoning process.
Chest X-Ray is the most commonly performed diagnostic imaging procedure worldwide, widely regarded as a cornerstone of modern radiology  (Johnson et al. 2019). However, interpreting CXRs demands careful multi-structure assessment that is time-consuming and expertise-intensive (Bahl, Ramzan, and Maraj 2020) In response, researchers have introduced a suite of task-specific models for classification (Rajpurkar, Irvin, and Zhu 2017), segmentation (Ma et al. 2024), and report generation (Tanno and Barrett 2024; Chambon and Delbrouck 2024), etc. While specialized AI tools for tasks like classification (Rajpurkar, Irvin, and Zhu 2017), segmentation (Ma et al. 2024) or report generation (Tanno and Barrett 2024; Chambon and Delbrouck 2024) have shown promise in improving turnaround time and diagnostic consistency (Baltruschat and Steinmeister 2021; Ahn 2022; Pham 2022; Shin 2023), their narrow specialization hinder their use in complex clinical reasoning scenarios (Erdal 2023; Fallahpour 2024).
Large-scale foundation models (FMs) in recent years like GPT-4o (OpenAI et al. 2024), LLaVA-Med (Li and Wong 2024), and CheXagent (Chen 2024) offer a more unified approach by integrating visual and textual reasoning. However, these monolithic systems often hallucinate (Eriksen, Möller, and Ryg 2024), lack domain-specific robustness (Chen 2024), and operate as uninterpretable "black boxes", making them unsuitable for high-stakes medical deployment.
Motivated by the need for more reliable, generalized, and autonomous solutions, recent efforts have explored multi-agent medical AI systems that coordinate domain-specific tools utilizing the capability of large language models (LLMs) and vision language models (VLMs). Recent progress in general-purpose agent systems (Li and Hammoud 2023; Wu et al. 2024; Zhuge 2024) demonstrate the potential of collaborative LLM agents to outperform single-agent baselines through structured communication and role specialization (Du et al. 2023; Liang 2023). Despite these advances, most systems rely on manually-defined and rigid workflows (Qian 2024; Zhang 2024b), which cannot adapt to the varying complexity of clinical queries and are computationally inefficient.
To address these challenges, recent methods have aimed to automate the design of multi-agent workflows. Works such as DsPy (Khattab 2023) and EvoPrompt (Guo 2023) optimize prompts, while G-Designer (Zhang 2024a) and AutoAgents (Chen et al. 2024a) refine inter-agent communication and profiling strategies. In the medical domain, MedRAX (Fallahpour et al. 2025) exemplifies this direction by orchestrating multiple CXR tools via ReAct-style prompting (Yao et al. 2023), achieving improved accuracy over end-to-end models. However, these methods largely rely on black-box LLMs for the decision-making of invoking agents, leaving the concerns regarding trustworthiness and safety risks as open questions.
The most recent advance, agentic supernets like MaAS (Zhang 2025), introduced a paradigm shift by learning a distribution over possible workflows, enabling adaptive, cost-aware reasoning. However, this approach has two fundamental flaws for medical applications. First, it is designed for text-only reasoning and lacks multimodal integration, which is inherently a core requirement in clinical reasoning. Second, while its textual gradient mechanism enables workflow optimization, it operates implicitly within the LLM’s internal prompt space during multi-turn conversations, providing limited true interpretability and traceability for end-users.
These challenges highlight a critical need for a medical agentic system that is not only multimodal and truly interpretable, but also adaptive and efficient. To this end, we propose PASS (Probabilistic Agentic Supernet Sampling). To the best of our knowledge, PASS is the first framework for interpretable and adaptive CXR reasoning via multimodal agentic workflow sampling. Given a CXR image and a complex free-form clinical reasoning task, PASS manages an evolving contextual memory, operates over a directed acyclic graph consisting of multiple specialized medical agent containers (i.e., agentic supernet), and adaptively samples layer-wise tool sequences from the graph. Crucially, we design a Controller module to learn the task-conditioned continuous distribution over the supernet, yielding decision paths annotated with interpretable probabilities. This provides transparent trajectories for post-hoc audits, directly enhancing medical AI safety. We design a principled three-stage regimen for the training of PASS: (1) expert knowledge-guided warm-up aligns tool usage with clinical best practices; (2) contrastive path-ranking sharpens ordering preferences among tool sequences; and (3) cost-aware reinforcement learning trains the controller to learn the optimized accuracy-cost Pareto frontier with an early-exit mechanism.
To systematically evaluate such agentic systems, we introduce ChestAgentBench-E (CAB-E), a challenging new benchmark comprising 2,550 comprehensive and safety-critical CXR reasoning cases annotated with free-form QA pairs, image inputs, and queries that highly demand complex rationales. CAB-E expands the scope of prior evaluations (Fallahpour et al. 2025; Liu et al. 2021), emphasizing multi-step and clinically grounded queries that require adaptive tool orchestration. We will make CAB-E publicly available upon publication.
Our key contributions can be summarized as follows:
  • We propose PASS, the first framework to instantiate a probabilistic agentic supernet for multimodal clinical reasoning, enabling an interpretable, adaptive, and cost-aware agentic system.
  • We design a principled three-stage training strategy including expert knowledge warm-up, contrastive path ranking, and cost-aware reinforcement learning.
  • We introduce CAB-E, a comprehensive public benchmark to evaluate complex and safety-critical agentic reasoning for CXR.
  • Extensive experiments show that PASS outperforms strong baselines among various benchmarks in accuracy, while maintaining the balanced computational cost and interpretable workflows.
PASS represents a paradigm shift for building trustworthy, adaptive, and efficient agentic systems for multimodal medicine reasoning.
2 Methodology
In this paper, we propose a probabilistic framework for PASS that interprets workflow construction as a latent decision-making process governed by a multimodal generative policy. In this section, we first formulate a probabilistic controller over tool trajectories and answers, and derive a cost-aware objective grounded in expected utility maximization. We then introduce the architecture and parameterization of the controller $\pi_{\theta}$, followed by a theoretically motivated multi-phase training algorithm that combines expert knowledge warm-up, contrastive path ranking and cost-aware reinforcement learning.
2.1 Preliminary
Problem formulation and notations.
Let $\mathcal{Q}=\{(q_{i},I_{i},C_{i})\}_{i=1}^{N}$ be a collection of multimodal diagnostic queries, where $q_{i}\!\in\!\mathcal{T}$ is a free-form text question, $I_{i}\!\in\!\mathbb{R}^{H\times W\times 3}$ is a chest X-ray image and $C_{i}\!\in\!\mathcal{C}$ denotes personalized contextual memory, including summarized information like structured demographic factors, clinical results, and previous analysis outputs. PASS answers $q_{i}$ by sampling a workflow $\tau$ over a directed acyclic multi-container graph and executing the tools in corresponding containers in the selected sequence. We frame workflow generation as sampling from a probability distribution $\pi_{\theta}$ based on multimodal evidence:
$\tau\sim\pi_{\theta}(\,\cdot\mid q,I,C),\quad\hat{a}=\textsc{Execute}(\tau)$ (1)
where the workflow $\tau=(a_{1},a_{2},\ldots,a_{T})$ is the trajectory of the actions and $\hat{a}$ is a free-form answer (e.g., finding, measurement, report section, etc.) returned to the clinician. PASS must simultaneously maximize diagnostic utility $\mathcal{U}$ and minimize a composite cost $\mathcal{L}$ capturing latency, token usage and privacy risk. The hyperparameter $\lambda$, configured by the user or deploying institution, controls the trade-off between performance and operational constraints. Under the above settings, the goal of our model could be stated as:
$\max_{\theta}\;\mathbb{E}_{\begin{subarray}{c}(q,I,C)\sim\mathcal{Q}\\ \tau\sim\pi_{\theta}\end{subarray}}\Bigl{[}\mathcal{U}(\hat{a},a^{\star})-\lambda\cdot\mathcal{L}(\tau)\Bigr{]}$ (2)
Agentic supernet.
Supernet $\mathcal{G}=(\mathcal{V},\mathcal{E})$ contains agent containers as nodes and legal tool invocations as edges. Each container $v\!\in\!\mathcal{V}$ is typed by one of Segmentation, Classify, Grounding, Report,  VQAnalyze, GuidelineLookup and MKG. The container $v$ also stores a mutable set of tool models $T_{v}=\{t_{v,1},\dots\}$ that share identical I/O signature but may differ in backbone architecture, patch size or training epoch. The detailed tool model descriptions are in the Appendix, Edges $e=(v\!\rightarrow\!v^{\prime})\!\in\!\mathcal{E}$ are labeled with a routing policy $\rho_{e}$ specifying which fields of the current memory are forwarded to the next container.
Formal Interface.
Every container $v$ adheres to a unified formal interface, defining its input $x_{v}$ and output $y_{v}$ as:
$\left\{\begin{aligned} x_{v}&=\bigl{(}q^{(\mathrm{sub})},\,I^{(\mathrm{roi})},\,C^{(\mathrm{sub})},\,\eta\bigr{)}\\ y_{v}&=\bigl{(}\rho_{v},\,\ell_{v},\,\kappa_{v}\bigr{)}\end{aligned}\right$. (3)
where $\rho_{v}$ is the primary multimodal payload. This payload can be a JSON object for structured data (e.g., {text, bbox, prob}) or an image tensor for visual data (e.g., a segmentation mask). The input $x_{v}$ consists of a textual sub-query $q^{(\mathrm{sub})}$, an optional region-of-interest image tensor $I^{(\mathrm{roi})}$, a relevant slice of personalized contexual memory $C^{(\mathrm{sub})}$, and tool-specific hyperparameters $\eta$. Finally, $\ell_{v}$ and $\kappa_{v}$ are the measured latency and token costs, respectively, used in Eq. (2). In our design, every container adheres to this interface specification, thus enabling plug-and-play maintenance.
Action space.
The space spanned by legal actions at state $s_{t}$ is defined as:
$\mathcal{A}(s_{t})=\bigl{\{}(v,T_{v,k})\mid(v_{t}\!\rightarrow\!v)\in\mathcal{E}\bigr{\}}\cup\{\textsc{EarlyExit}\}$
where $(v,T_{v,k})$ denotes executing tool $T_{v,k}$ inside container $v$. Sampling the EarlyExit action, a special action that halts the execution trajectory early to conserve resources, thus initiating answer synthesis in advance.
Algorithm 1 PASS: Training Procedure
0: Expert demonstrations $\mathcal{D}_{\mathrm{exp}}$, unlabeled data $\mathcal{D}_{\mathrm{ul}}$, supernet $\mathcal{G}$, state encoder $\psi$, policy $\pi_{\theta}$, answer generator $p_{\phi}$, heuristic reward $R_{h}$, cost weights $\lambda$, entropy weight $\gamma$. Procedure 1
1:# Phase I: Expert Knowledge Warm-up
2:for $(s,a^{\star})\in\mathcal{D}_{\mathrm{exp}}$ do
3:  $\theta\leftarrow\theta-\eta_{1}\nabla_{\theta}\bigl{(}-\log\pi_{\theta}(a^{\star}\mid s)\bigr{)}$
4:end for
5:# Phase II: Contrastive Path Ranking
6:for $(q,I,C)\in\mathcal{D}_{\mathrm{ul}}$ do
7:  $\{\tau_{k}\}_{k=1}^{K}\sim\pi_{\theta}(\cdot\mid q,I,C)$
8:  $p(\tau_{k})\leftarrow\frac{\exp(R_{h}(\tau_{k})/\alpha_{\mathrm{cpr}})}{\sum_{j=1}^{K}\exp(R_{h}(\tau_{j})/\alpha_{\mathrm{cpr}})}$
9:  $\mathcal{L}_{\textsc{CPR}}\leftarrow-\sum_{k=1}^{K}p(\tau_{k})\log\pi_{\theta}(\tau_{k})$
10:  Update $\theta$ using $\nabla_{\theta}\mathcal{L}_{\textsc{CPR}}$
11:end for
12:# Phase III: Cost-aware Reinforcement Learning
13:for $n=1$ to $N_{\mathrm{RL}}$ do
14:  $\tau\sim\pi_{\theta}(\cdot\mid q,I,C)\in\mathcal{D}_{\mathrm{ul}}$
15:  $\hat{a}\sim p_{\phi}(\cdot\mid\tau,q,I,C)$
[…]
17:  $\theta\leftarrow\theta+\eta_{3}R(\tau)\nabla_{\theta}\log\pi_{\theta}(\tau)$
18:end for
Algorithm 2 PASS: Inference
0: Policy $\pi_{\theta}$, generator $p_{\phi}$, summarizer $\mathcal{S}$, state encoder $\psi$, supernet $\mathcal{G}$, max steps $T_{\max}$. Procedure 2
[…]
2:for $t=1$ to $T_{\max}$ do
3:  $a_{t}\sim\pi_{\theta}(\cdot\mid\psi(q,I,C,M))$
4:  if $a_{t}=\textsc{EarlyExit}$ then
5:   break
6:  end if
7:  $\rho_{t}\leftarrow\textsc{ExecuteTool}(a_{t})$
8:  $M\leftarrow M\cup\mathcal{S}(\rho_{t})$
9:  $\tau\leftarrow\tau\cdot a_{t}$
10:end for
[…]
12:Return $(\hat{a},\tau)$// Return final answer and full workflow
2.2 Multi-agent Workflows
PASS models diagnostic reasoning as structured decision-making in a latent space of tool-based workflows. Given an input triplet $(q,I,C)$, the agent sequentially builds a trajectory $\tau=(a_{1},a_{2},\ldots,a_{T})$ by sampling actions $a_{t}\in\mathcal{A}(s_{t})$, where $s_{t}$ is the multimodal reasoning state at step $t$. The agent’s final output is a multimodal package, consisting of a final textual answer $\hat{a}\in\mathcal{T}$ and any visual artifacts (e.g., annotated images) produced during the workflow $\tau$.
The core of PASS is the workflow policy $\pi_{\theta}(a_{t}\mid s_{t})$, which we aim to learn. This policy, combined with a fixed answer synthesis module $p_{\phi}$, defines the full generative process for the textual answer $\hat{a}$:
$p_{\theta}(\tau,\hat{a}\mid q,I,C)$ (4)
$=$ $\underbrace{p_{\phi}(\hat{a}\mid\tau,q,I,C)}_{\text{Answer generator}}\cdot\underbrace{\prod_{t=1}^{T}\pi_{\theta}(a_{t}\mid s_{t})}_{\text{Workflow policy}}$
where $p_{\phi}$ is a frozen synthesis module (e.g., a large language model) responsible for generating the final text answer $\hat{a}$ based on the evidence gathered in $\tau$. All learning is concentrated in the policy parameters $\theta$, ensuring improvement stems from discovering better workflow decisions, not from fine-tuning the generator. This decomposition makes two key assumptions: (i) the tool sampling is a Markov process over the state space $\mathcal{S}$, and (ii) the final textual answer is conditionally independent of the internal policy decisions, given the full trajectory $\tau$.
Policy-induced answer distribution.
By virtue of marginalizing out the latent tool trajectory $\tau$, we obtain the model-implied distribution over answers:
$p_{\theta}(\hat{a}\mid q,I,C)$ (5)
$=$ $\sum_{\tau\in\mathcal{T}(q,I,C)}\pi_{\theta}(\tau\mid q,I,C)\cdot p_{\phi}(\hat{a}\mid\tau,q,I,C)$
in which $\pi_{\theta}(\tau\mid q,I,C)=\prod_{t=1}^{|\tau|}\pi_{\theta}(a_{t}\mid s_{t})$ and $\mathcal{T}(q,I,C)$ is the set of legal trajectories under $\mathcal{G}$ from initial state $s_{0}$. Although this marginal distribution is intractable to compute exactly due to the combinatorial size of $\mathcal{T}$, it can be approximated with Monte Carlo sampling, which we exploit both for training and for uncertainty estimation.
Expected utility and cost regularization.
Given a ground-truth answer $a^{\star}$ and a reward function $\mathcal{U}(\hat{a},a^{\star})$ measuring the clinical utility of the predicted answer, our goal is to maximize the expected utility of our policy. This objective must also be balanced against the cost of the workflows it generates. Formally, the goal is to find optimal parameters $\theta$ for the policy $\pi_{\theta}$:
$\max_{\theta}\;\;\mathbb{E}_{(q,I,C)\sim\mathcal{Q}}\Bigl{[}\mathbb{E}_{\hat{a}\sim p_{\theta}(\cdot\mid q,I,C)}\mathcal{U}(\hat{a},a^{\star})$
$-\lambda\cdot\mathbb{E}_{\tau\sim\pi_{\theta}(\cdot\mid q,I,C)}\mathcal{L}(\tau)\Bigr{]}$ (6)
This formulation can be viewed as a constrained variational inference problem over the latent workflow $\tau$ with an amortized inference network $\pi_{\theta}$.
Uncertainty-aware generation.
The posterior entropy of the answer distribution, $H_{\theta}(\hat{a}\mid q,I,C)=-\mathbb{E}_{\hat{a}}\log p_{\theta}(\hat{a}\mid q,I,C)$, can be utilized to quantify the epistemic uncertainty of the model. Since the answer generator $p_{\phi}$ is frozen, this entropy is solely induced by the sampling variability in the workflow trajectory $\tau\sim\pi_{\theta}$. In practice, we estimate $H_{\theta}$ via Monte Carlo rollouts of the policy and use it both as a proxy for answer confidence and as a regulariser during policy learning (Sec. 2.4) to discourage high-entropy outputs in high-risk settings.
2.3 Controller Architecture
The controller $\pi_{\theta}(a_{t}\mid s_{t})$ is designed as a masked categorical distribution over permissible actions, with its parameters determined by a state encoder $\psi$ Its logits are produced by a policy network head that processes the state representation $h_{t}$. Let $s_{t}=(q,I,C,M_{t})$ denote the current multimodal state. The state encoder maps this input into a shared representation $h_{t}\in\mathbb{R}^{d}$:
$\begin{gathered}h_{t}=\psi(s_{t})=\mathrm{LN}(z_{t})\quad s.t.\\ z_{t}=W_{I}\cdot\xi(I)\;\|\;W_{Q}\cdot\zeta(q,C)\;\|\;W_{M}\cdot\mu(M_{t})\end{gathered}$ (7)
where $\xi(I)$ is a frozen ViT-B/16 image encoder with final-layer CLS token projected to $\mathbb{R}^{256}$, $\zeta(q,C)$ is a Sentence-BERT-style text encoder for $(q,C)$, projected to $\mathbb{R}^{128}$, $\mu(M_{t})$ encodes memory as a 2-layer Transformer over token summaries (max length 256), with pooled final hidden state $\in\mathbb{R}^{128}$, $\mathrm{LN}(\cdot)$ denotes layer normalization, and $\|$ denotes concatenation. The policy head is a feed-forward network with a single hidden layer and ReLU activation:
$\pi_{\theta}(a_{t}\mid s_{t})$ (8)
$=$ $\mathrm{Softmax}\left(\mathrm{mask}_{\mathcal{A}(s_{t})}\left[W_{2}\cdot\sigma(W_{1}h_{t})\right]\,/\,\alpha\right)$
where the legal-action mask $\mathrm{mask}_{\mathcal{A}(s_{t})}$ zeroes out infeasible transitions in the supernet $\mathcal{G}$ and $\alpha$ is a temperature parameter annealed during training from 2.0 to 0.8.
Personalized contexutal memory.
At step $t$, what the controller observes are stated as:
$s_{t}=\Bigl{(}q,I,C,M_{t}\Bigr{)},\quad M_{t}=\bigl{\{}(v_{j},\tilde{y}_{v_{j}})\bigr{\}}_{j=1}^{t-1}$
where the memory $M_{t}$ is a bounded-size first-in-first-summarized (FIFS) buffer. After each tool call, in order to save the computational cost, the JSON response $y_{v}$ is summarized to a compressed vector $\tilde{y}_{v}$ using a frozen language model prompted to function only as paraphrasing. These textual summaries are appended to a FIFO memory $M_{t}$ along with image outputs (if any). This personalized and evolving memory mechanism enables precise, in-context diagnosis in the wild.
Table 1: Performance across three radiology VQA benchmarks (mean $\pm$ standard deviation). Best and runner-up numbers are bold and underlined.
[…]
2.4 Three-Stage Training Procedure
We train the workflow policy $\pi_{\theta}$ to optimize the objective in Eq. (6) via a principled three-stage procedure. This curriculum-based approach progressively refines the policy, starting with strong expert supervision before moving to weaker preference signals and finally to direct reinforcement learning on the end-task reward. The three stages are detailed as follows. Each stage is grounded in a formal objective, allowing for stable and efficient training of $\pi_{\theta}$.
Phase I: Expert knowledge guided warm-up.
This initial phase uses imitation learning to bootstrap the policy. We construct a dataset of expert demonstrations, $\mathcal{D}_{\mathrm{exp}}$, not from scratch, but by using a more scalable, two-step process. First, we use a powerful foundation model (GPT-4o) to generate initial workflow sketches for a set of problems. Second, these sketches are then reviewed, corrected, and validated in a human-in-the-loop process by licensed radiologists. This “distill-and-refine” strategy yields a high-quality dataset of one-step decisions $\mathcal{D}_{\mathrm{exp}}=\{(s,a^{\star})\}$, where $a^{\star}$ is the expert-verified action for state $s$. We warm-start the policy by minimizing the KL divergence from the expert policy (i.e., behavior cloning):
$\mathcal{L}_{\textsc{BC}}=\mathbb{E}_{(s,a^{\star})\sim\mathcal{D}_{\mathrm{exp}}}\left[-\log\pi_{\theta}(a^{\star}\mid s)\right]$ (9)
This phase instills a strong prior in the policy, anchoring it in clinically valid reasoning patterns.
Phase II: Heuristic-guided contrastive path ranking.
Expert demonstrations are costly to acquire and cannot cover all scenarios. To generalize beyond $\mathcal{D}_{\mathrm{exp}}$, we introduce a weaker supervisory signal based on heuristic preferences for unlabeled data. For a given query, we sample $K$ candidate workflows $\{\tau_{k}\}_{k=1}^{K}$ from the current policy $\pi_{\theta}$. We then score each path using a heuristic reward function, $R_{h}(\tau_{k})$, which combines domain-specific priors such as clinical guideline compliance, anatomical coherence, and brevity. The policy is then updated using a contrastive loss (InfoNCE) that encourages it to assign higher probability to higher-scoring paths:
$\begin{split}\mathcal{L}_{\textsc{CPR}}=&\mathbb{E}_{\{\tau_{k}\}\sim\pi_{\theta}}\left[-\sum_{k=1}^{K}p(\tau_{k})\log\pi_{\theta}(\tau_{k})\right],\\ &\text{where }p(\tau_{k})=\frac{\exp(R_{h}(\tau_{k})/\alpha_{\mathrm{cpr}})}{\sum_{j=1}^{K}\exp(R_{h}(\tau_{j})/\alpha_{\mathrm{cpr}})}\end{split}$ (10)
where $\alpha_{\mathrm{cpr}}$ is a temperature hyperparameter. This phase teaches the policy to distinguish between good and bad reasoning structures, even without a ground-truth workflow.
Phase III: Cost-aware policy refinement.
In the final phase, we directly fine-tune the policy $\pi_{\theta}$ using reinforcement learning to maximize the expected end-task utility. To compute the reward for a generated workflow $\tau$, we first use the fixed answer generator $p_{\phi}$ to synthesize a textual answer, $\hat{a}\sim p_{\phi}(\cdot\mid\tau,q,I,C)$. We then define the reward for the trajectory as:
[…]
where H(a^)H($\hat${a}) is the entropy of generated answers, penalizing uncertainty. We then update the policy parameters $\theta$ using a reinforcement learning approach. The objective is to maximize the expected reward over all trajectories sampled from the policy:
$J(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}[R(\tau)]$ (12)
The gradient of this objective, $\nabla_{\theta}J(\theta)$, can be estimated using sampling like the reinforcement algorithm, often with a baseline to reduce variance. This final tuning step aligns the workflow generation directly with the ultimate goals of diagnostic accuracy and computational efficiency.
3 Experiments
We evaluate PASS across three radiology benchmarks of increasing complexity to assess four critical aspects of real-world deployment: clinical accuracy, language fidelity, computational efficiency, and safety. All experiments are conducted on a single NVIDIA H800 (80GB) GPU with access to OpenAI’s GPT-4o API for relevant baselines.
3.1 Experiment Setup
Benchmarks. We use the following evaluation suites, with more details decribed in the Appendix:
  • CAB-E: 2,550 multi-step chest X-ray (CXR) reasoning cases, including 500 safety-critical instances, of which construction details and statistics summary are attached in the Appendix. This benchmark targets free-form, multi-hop reasoning grounded in both image and patient context. The safety split emphasizes complex scenarios which require careful and transparent decision-making process, including life-threatening anatomical abnormalities, urgent systemic conditions, etc. We include more details in the Appendix. It will be made publicly available upon publication of the paper.
  • CAB-Standard (Fallahpour et al. 2025): A multiple-choice Chest Agent Benchmark (CAB) containing 2,500 diagnostic queries. CAB-Standard is constructed using the generation method proposed by Fallahpour et al.
  • SLAKE (Liu et al. 2021): A native free-form medical VQA benchmark with 6,437 image–question pairs, used to assess zero-shot generalization.
Metrics. On CAB-E, we report: Accuracy, LLM-as-a-Judge score (LLM-J.), BLEU, METROR, ROUGE-L, embedding similarity, and end-to-end latency. CAB-Standard is evaluated by accuracy and latency. SLAKE is evaluated by answer AUROC and latency. We also evaluate the hallucination rate on the safety-critical split of CAB-E. Detailed descriptions of the metrics can be found in the Appendix.
[…]
Implementation details.
We optimize the model using the AdamW algorithm, incorporating gradient clipping at 1.0 to ensure numerical stability, a weight decay of 0.01 to prevent overfitting, and a cosine learning rate schedule to facilitate smooth convergence. An entropy bonus of 0.01 is applied to encourage exploration and stabilize training. For RL updates, we employ forward-mode unrolling with a 5-step truncation to balance computational efficiency and gradient accuracy.
Table 2: Performance on 500 radiologist-verified safety-critical cases from CAB-E.
[…]
GPT-4o (zero-shot) 61.22 7.00
LLaVA-Med 87.75 2.00
MedRAX 89.79 1.60
PASS 93.50 1.60
[…]
Full PASS 90.16 -
   – EarlyExit 88.60 94.0
   – Path-Rank Pretraining 87.86 8.9
   – Expert-Guided Warm-up 88.89 9.5
Figure 2: Cost-Accuracy Pareto Frontier for PASS. Lower normalized cost and higher accuracy are preferred. Each point corresponds to a specific configuration of penalty weights, allowing flexible trade-offs during deployment. The arrow denotes the preferred direction (i.e. towards the top-left region), indicating simultaneously reduced cost and improved accuracy.
3.2 Main Results on CAB-E
Table 1 presents the results on CAB-E. PASS achieves an accuracy of 91.22, outperforming the strongest baseline MedRax (89.59) by +1.54, surpassing CheXagent by +7.55and LLaVA-Med by +4.26, demonstrating substantial improvement in diagnostic accuracy through probabilistic multi-tool reasoning. This suggests that adaptively sampled agentic trajectories, rather than static templates, offer superior coverage and reliability on diverse CXR cases.
PASS also achieves the highest LLM-J. score (84.28), BLEU (8.51), METEOR (33.21) and ROUGE-L score (31.49). This indicates that the generated impressions align better with ground truth clinical solutions, validating the controller’s ability to coordinate image grounding and textual fluency across multi-hop tool outputs.
Latency and cost-performance analysis.
The Pareto frontier in Figure 2 visualizes our framework’s design philosophy. While PASS exhibits higher latency than single-pass models like LLaVA-Med (Table 1), this is a direct and strategic trade-off for its superior accuracy, driven by a more comprehensive reasoning process. This is exemplified by PASS-High (91.22% Acc.), incurring the maximum normalized cost (1.0) to generate a transparent, auditable workflow essential for high-stakes clinical safety, which is a safeguard absent in opaque ’black-box’ models. More importantly, unlike baselines which represent fixed performance-cost points, PASS offers a spectrum of operational choices. This enables the selection of a balanced configuration like PASS-Medium (Normalized cost: 0.708, Acc: 88.0%) for screening or the high-fidelity PASS-High for complex diagnostics, facilitating strategic resource allocation in real-world scenarios.
3.3 Safety-Critical Subset Evaluation
On this high-risk subset, PASS achieves a remarkable accuracy of 93.50%, surpassing MedRAX by 3.71% and LLaVA-Med by 5.75%. Notably, PASS matches MedRAX with the lowest hallucination rate of 1.60%, a significant improvement compared to GPT-4o’s 7.00%, highlighting its robustness in minimizing errors in safety-critical medical cases.. These results highlight PASS’s exceptional reliability in safety-critical medical scenarios.
3.4 Cost–Accuracy Trade-off
Figure 2 illustrates the empirical Pareto frontier of PASS by varying the penalty weights $\lambda$. The x-axis denotes normalized inference cost (relative to GPT-4o), and the y-axis reports accuracy. PASS enables flexible trade-offs: increasing $\lambda$ reduces cost substantially with minimal accuracy loss. For example, $\lambda=0.03$ lowers cost by 18% (from 1.00 to 0.827) while retaining 88.1% accuracy. The highest accuracy (91.2%) is achieved at $\lambda=0.003$ with modest cost, and the best cost-adjusted accuracy arises at $\lambda=0.3$, which halves the cost while maintaining strong performance. This demonstrates that PASS learns a well-structured frontier, enabling deployment-time adaptation based on resource constraints.
3.5 Ablation Study
Ablation results (Table 3) confirm critical design choices: Removing early-exit causes a significant accuracy drop (88.60 vs 90.16) and a 94% cost increase. Removing path-rank pretraining and warm-up also demonstrates their role in convergence acceleration and performance improvements.
4 Related Works
Tool-augmented LLMs.
Tool use in LLMs has evolved from basic augmentation (Schick et al. 2023; Yao et al. 2023; Feng et al. 2025b) to modular agent frameworks (Wu et al. 2024; Li and Hammoud 2023; Chen et al. 2024b; Zhuge et al. 2024) with specialized roles and communication. Yet, most rely on static or handcrafted workflows, lacking adaptability and efficiency in real-world deployment. While recent efforts explore reinforcement learning for tool strategies (Feng et al. 2025a), few support uncertainty modeling or dynamic inference structures.
Autonomous agent workflows.
Recognizing the limitations of fixed pipelines, a new wave of research seeks to automate agentic system design. Prompt optimization (Khattab 2023; Guo 2023), inter-agent communication tuning (Zhang 2024a), and modular profiling (Chen et al. 2024a) form key directions. Notably, MaAS (Zhang et al. 2025) introduces a paradigm shift: instead of selecting a single best workflow, it learns a distribution over agentic architectures, enabling query-dependent inference. However, existing approaches are largely limited to text-only domains and provide limited interpretability, being particularly problematic in high-stakes applications such as medicine.
Multimodal reasoning in medical AI.
Multimodal foundation models (e.g., GPT-4V (Nori 2023), LLaVA-Med (Li and Wong 2024), CheXagent (Chen 2024)) promise unified vision-language understanding and have shown zero-shot capabilities across radiological tasks. Still, they often hallucinate (Eriksen, Möller, and Ryg 2024), lack task specificity (Chen 2024), and remain opaque. Domain-specific systems like MedRAX (Fallahpour et al. 2025) and MDAgents (Kim et al. 2024) attempt to integrate medical tools with LLMs via ReAct-style (Yao et al. 2023) prompting, offering partial medical multimodal reasoning capabilities. Yet, their decision-making still largely rely on black-box LLMs, hindering real-world application due to critical concerns about trust and potential risks.
Resource-aware learning.
To achieve both accuracy and efficiency, recent studies have explored adaptive computation via early-exit transformers (Kang et al. 2024), budgeted RL (Carrara et al. 2019), and adaptive computation graphs (Graves 2016). GFlowNets (Bengio et al. 2021) offer a general framework for sampling diverse, high-reward trajectories in structured spaces, and have been applied in molecular design (Chen et al. 2024a) and agentic planning (Zhang et al. 2025).
Safety and interpretability in clinical deployment.
Clinical settings demand more than performance: they require transparency, controllability, and regulatory compliance (Lundervold and Lundervold 2019). Beyond saliency-based explanations, methods like MedCoT (Liu et al. 2024) and BoxMed-RL (Jing et al. 2025) leverage chain-of-thought or RL-enhanced generation to increase reliability. PASS extends these efforts with per-step, probability-annotated execution traces and interpretable early exits, allowing for post-hoc audits and fine-grained trust calibration, which are crucial features for safe medical AI deployment.
5 Conclusion
In this paper, we introduced PASS, the first multimodal framework to address the critical challenges of interpretability, adaptability, and efficiency in complex chest X-ray reasoning. Existing agentic systems are often limited by their black-box nature, poor integration of multimodal data, and rigid, inefficient workflows. PASS overcomes these limitations by leveraging a probabilistic controller to adaptively sample workflows from a multi-tool supernet, yielding decision paths annotated with transparent probabilities that are crucial for clinical trust and post-hoc audits. Our novel three-stage training regimen which includes expert knowledge warm-up, contrastive path-ranking, and cost-aware reinforcement learning to optimize the performance-cost trade-off, balancing diagnostic accuracy with computational cost via a dynamic early-exit mechanism. Through extensive experiments on our newly curated CAB-E benchmark and other public datasets, we have demonstrated that PASS not only achieves superior accuracy over strong baselines but also provides interpretable and efficient reasoning. Ultimately, PASS represents a significant paradigm shift, paving the way for the next generation of trustworthy, adaptive, and resource-aware agentic systems for high-stakes multimodal medical applications.
Limitations
PASS deliberately uses a fixed container set to ensure clinical safety and interpretability as a strategic trade-off between safety and flexibility. Future works will scale this robust foundation by expanding the supernet to new imaging types like MRI or CT, and enriching its agentic containers and tools, thereby further enhancing the diagnostic utility and adaptability of PASS.